# Magic Gems: A Polyhedral Framework for Magic Squares Connecting Combinatorics, Geometry, and Linear Algebra

### v3

@@ 1 Introduction
Magic squares have fascinated mathematicians, artists, and mystics for millennia [11]. The oldest known example, the Lo Shu square, appears in Chinese legend dating back to 2800 BCE and remains the unique (up to symmetry) $3\times 3$ magic square using the integers 11 through 99. Despite this ancient pedigree and centuries of intensive study [2], magic squares continue to yield new mathematical insights, with recent work connecting them to algebraic structures, eigenvalue problems, and statistical mechanics.
@@
In this paper, we develop a geometric perspective on magic squares by representing them as three-dimensional polyhedra, which we call Magic Gems. The construction arises from a natural physical analogy: imagine placing checkers on an $n\times n$ board, where the height of each stack corresponds to the number occupying that cell. This creates a “topography” over the grid, and centering this configuration in three-dimensional space yields a point cloud whose convex hull defines the Magic Gem. This geometric approach complements prior work by Loly and collaborators on the physical properties of magic squares [6, 7, 9].
@@ 1.2 Paper Organization
The remainder of this paper proceeds as follows. Section 2 reviews classical results on magic squares and establishes notation for the geometric and statistical concepts we employ. Section 3 presents the Magic Gem construction, proves the Complete Energy Characterization Theorem, introduces the low-mode relaxation, and develops the perturbation analysis framework. Section 4 describes our computational experiments, validating theoretical predictions and exploring the structure of the arrangement space for $n=3$, 44, and 55. Section 5 interprets our findings and discusses connections to related areas. Finally, Section 6 summarizes our contributions and outlines directions for future research.
@@ 2.1 Magic Squares: Classical Results
The enumeration of magic squares reveals a rapid growth in complexity with the order $n$ [12]. For $n=3$, there exist exactly eight magic squares, all related by rotations and reflections, yielding a single essentially different magic square—the Lo Shu. The situation changes dramatically for larger orders: $n=4$ admits 7,040 magic squares forming 880 equivalence classes under the dihedral group, while $n=5$ yields approximately 275 million magic squares comprising roughly 34 million equivalence classes. This explosive growth poses significant challenges for enumeration and classification, motivating the search for structural insights that transcend explicit construction.
@@ 2.4 Covariance and the Moment of Inertia
The eigenvalues of II, called the principal moments of inertia, characterize the rotational properties of the point configuration and are invariant under rigid rotations. For equal unit masses, the inertia tensor and covariance matrix are closely related, differing by a trace term and sign conventions.
@@ 3.1 From Magic Square to Point Cloud
with $n=3$ and center value 55, the nine points are
@@ 4.4 Cross-Order Comparison and Scaling Analysis
The zero-covariance theorem holds universally across orders, but the geometric and energy landscape properties evolve systematically with $n$. Figure 5 presents a unified view of Magic Gems across orders 33, 44, and 55, along with key scaling metrics.
@@
Several systematic patterns emerge. The convex hull volume grows rapidly with $n$, from 18.0 for $n=3$ to 276.7 for $n=5$, reflecting both the expanding $z$-value range (4-4 to 44 for $n=3$, versus 12-12 to $12$ for $n=5$) and the increasing number of vertices. The fraction of points lying on the hull boundary decreases from $8/9\approx 0.89$ for $n=3$ to roughly $0.64$ for $n=5$, as the hull interior becomes more populated.
@@ 5.1.2 Physical and Algebraic Interpretations
From a physical perspective, the zero-covariance property implies that the Magic Gem, viewed as a rigid body with masses at its vertices, has a simplified inertia tensor. The vanishing of $I_{xz}=-n^{2}\mathrm{Cov}(X,Z)$ and $I_{yz}=-n^{2}\mathrm{Cov}(Y,Z)$ means that the xzxz and yzyz products of inertia are zero. This implies the $z$-axis is aligned with a principal axis of the inertia tensor, simplifying the rotational dynamics about this direction.
@@ 5.3 Scaling Laws and Landscape Evolution
Across over 460 million sampled arrangements—including exhaustive enumeration for $n=3$, 60 million samples for $n=4$, and 400 million for n=5n=5—no zero-energy configurations of the low-mode energy $E_{\mathrm{low}}$ were discovered by random sampling. However, this does not imply that magic squares are the only zeros of $E_{\mathrm{low}}$ for $n\geq 4$; as shown in Remark 3.10, non-magic zeros exist but are too rare for random sampling to detect. The low-mode and complete energies measure different aspects of balance, and only the complete energy $E_{\mathrm{full}}$ (Theorem 3.17) provides a rigorous characterization for all orders. The probability of discovering a magic square by random sampling decreases super-exponentially with $n$, from roughly 11 in 45,00045{,}000 for $n=3$ to about $2\times 10^{-17}$ for $n=5$ (roughly one in $5\times 10^{16}$). This quantifies the “needle in a haystack” intuition with concrete numbers.
@@ 5.6.1 Physical Interpretations
Loly and collaborators have extensively developed the physical interpretation of magic squares [6, 7, 8, 9]. In seminal work, Loly [7] showed that when the entries of a magic square are treated as point masses, the moment of inertia tensor about the center exhibits remarkable invariance properties: all three principal moments are equal, making the configuration equivalent to a spherical top. This invariance is a direct consequence of the balanced row, column, and diagonal sums.
@@ 5.6.2 Spectral Properties
Previous work has also characterized magic squares through their eigenvalue structure, showing that the matrix of a magic square has predictable spectral properties [6, 9, 10, 5]. Connections to permutation matrices have also been explored [4]. Our covariance characterization complements these algebraic approaches: while eigenvalues describe the linear operator defined by the matrix, covariances describe the statistical structure of the associated point cloud. Both perspectives capture aspects of the “balance” that defines magic squares.

### v4

@@ 1 Introduction
Magic squares have fascinated mathematicians, artists, and mystics for millennia [13], appearing famously in Albrecht Dürer’s 1514 engraving Melencolia I [7]. The oldest known example, the Lo Shu square, appears in Chinese legend dating back to 2800 BCE and remains the unique (up to symmetry) $3\times 3$ magic square using the integers $1$ through $9$. Despite this ancient pedigree and centuries of intensive study [3], magic squares continue to yield new mathematical insights, with recent work connecting them to algebraic structures, eigenvalue problems, and statistical mechanics.
@@
In this paper, we develop a geometric perspective on magic squares by representing them as three-dimensional polyhedra, which we call Magic Gems. The construction arises from a natural physical analogy: imagine placing checkers on an $n\times n$ board, where the height of each stack corresponds to the number occupying that cell. This creates a “topography” over the grid, and centering this configuration in three-dimensional space yields a point cloud whose convex hull defines the Magic Gem. This geometric approach complements prior work by Loly and collaborators on the physical properties of magic squares [8, 10].
@@ 1.2 Paper Organization
The remainder of this paper proceeds as follows. Section 2 reviews classical results on magic squares and establishes notation for the geometric and statistical concepts we employ. Section 3 presents the Magic Gem construction, proves the Complete Energy Characterization Theorem, introduces the low-mode relaxation, and develops the perturbation analysis framework. Section 4 describes our computational experiments, validating theoretical predictions and exploring the structure of the arrangement space for $n=3$, 44, and $5$. Section 5 interprets our findings and discusses connections to related areas. Finally, Section 6 summarizes our contributions and outlines directions for future research.
@@ 2.1 Magic Squares: Classical Results
The enumeration of magic squares reveals a rapid growth in complexity with the order $n$ [14, 5]. For $n=3$, there exist exactly eight magic squares, all related by rotations and reflections, yielding a single essentially different magic square—the Lo Shu. The situation changes dramatically for larger orders: $n=4$ admits 7,040 magic squares forming 880 equivalence classes under the dihedral group, while $n=5$ yields approximately 275 million magic squares comprising roughly 34 million equivalence classes. This explosive growth poses significant challenges for enumeration and classification, motivating the search for structural insights that transcend explicit construction.
@@ 2.4 Covariance and the Moment of Inertia
The eigenvalues of $I$, called the principal moments of inertia, characterize the rotational properties of the point configuration and are invariant under rigid rotations. For equal unit masses, the inertia tensor and covariance matrix are closely related, differing by a trace term and sign conventions.
@@ 3.1 From Magic Square to Point Cloud
with $n=3$ and center value $5$, the nine points are
@@ 4.4 Cross-Order Comparison and Scaling Analysis
The zero-covariance theorem holds universally across orders, but the geometric and energy landscape properties evolve systematically with $n$. Figure 5 presents a unified view of Magic Gems across orders $3$, 44, and $5$, along with key scaling metrics.
@@
Several systematic patterns emerge. The convex hull volume grows rapidly with $n$, from 18.0 for $n=3$ to 276.7 for $n=5$, reflecting both the expanding $z$-value range (4-4 to $4$ for $n=3$, versus 12-12 to $12$ for $n=5$) and the increasing number of vertices. The fraction of points lying on the hull boundary decreases from $8/9\approx 0.89$ for $n=3$ to roughly $0.64$ for $n=5$, as the hull interior becomes more populated.
@@ 5.1.2 Physical and Algebraic Interpretations
From a physical perspective, the zero-covariance property implies that the Magic Gem, viewed as a rigid body with masses at its vertices, has a simplified inertia tensor. The vanishing of $I_{xz}=-n^{2}\mathrm{Cov}(X,Z)$ and $I_{yz}=-n^{2}\mathrm{Cov}(Y,Z)$ means that the $xz$ and $yz$ products of inertia are zero. This implies the $z$-axis is aligned with a principal axis of the inertia tensor, simplifying the rotational dynamics about this direction.
@@ 5.3 Scaling Laws and Landscape Evolution
Across over 460 million sampled arrangements—including exhaustive enumeration for $n=3$, 60 million samples for $n=4$, and 400 million for n=5n=5—no zero-energy configurations of the low-mode energy $E_{\mathrm{low}}$ were discovered by random sampling. However, this does not imply that magic squares are the only zeros of $E_{\mathrm{low}}$ for $n\geq 4$; as shown in Remark 3.10, non-magic zeros exist but are too rare for random sampling to detect. The low-mode and complete energies measure different aspects of balance, and only the complete energy $E_{\mathrm{full}}$ (Theorem 3.17) provides a rigorous characterization for all orders. The probability of discovering a magic square by random sampling decreases super-exponentially with $n$, from roughly $1$ in 45,00045{,}000 for $n=3$ to about $2\times 10^{-17}$ for $n=5$ (roughly one in $5\times 10^{16}$). This quantifies the “needle in a haystack” intuition with concrete numbers.
@@ 5.6.1 Physical Interpretations
Loly and collaborators have extensively developed the physical interpretation of magic squares [8, 9, 10]. In seminal work, Loly [8] showed that when the entries of a magic square are treated as point masses, the moment of inertia tensor about the center exhibits remarkable invariance properties: all three principal moments are equal, making the configuration equivalent to a spherical top. This invariance is a direct consequence of the balanced row, column, and diagonal sums.
@@ 5.6.2 Spectral Properties
Previous work has also characterized magic squares through their eigenvalue structure, showing that the matrix of a magic square has predictable spectral properties [10, 11, 12]. Connections to permutation matrices and polytope enumeration have also been explored [6, 2]. Our covariance characterization complements these algebraic approaches: while eigenvalues describe the linear operator defined by the matrix, covariances describe the statistical structure of the associated point cloud. Both perspectives capture aspects of the “balance” that defines magic squares.